OpenStudy (anonymous):

Evaluate improper integral dt/(t^2-2t)

OpenStudy (anonymous):

dt/(t^2-2t)=[-1/(2t)+1/(2{t-2})]dt = -1/2 ln (t)+1/2 ln (t-2) +c